by using jean number and average energy derive planks law of radiation
1
Expert's answer
2020-01-27T09:34:02-0500

The average energy E of an oscillator is connect with density σ of radiation by formula

"σ(υ, T)=\frac{8πυ^2}{c^3} \times {E} (1)"

where υ is oscillator frequency, c is the speed of light


The average energy is given by formula

"E=\frac{hυ}{\exp{\frac{hυ}{kT}}-1} (2)"

where h is the Planck’s constant, T is the temperature, k is the Boltzman’s constant


We put (2) in (1)

"σ(υ, T)=\frac{\frac{8πhυ^3}{c^3}}{\exp{\frac{hυ}{kT}}-1} (3)"

It is a Planck’s law of radiation.